OpenStudy (anonymous):

WILL MEDAL. a collection of books is set up into 34 rows. The first row has 96 books, and each row after that one has 2 less books then the preceding row. How many books are in the 17th row?

OpenStudy (dtan5457):

I would go with an equation like 96-2(y-1)=amount of books y=rows it's y-1 because 2 rows, should only have 2 less, not 4.

OpenStudy (triciaal):

maybe binomial theorem

OpenStudy (anonymous):

i think it is either 164 or 255

OpenStudy (dtan5457):

@sar12389 There is no way the 17th row will have more books than the first.

OpenStudy (dtan5457):

96-2(y-1)=amount of books y=rows plug into that equation^

OpenStudy (dtan5457):

y=17 in this case

OpenStudy (anonymous):

i read the question wrong then. The answer would be 64

OpenStudy (anonymous):

the possible answers are 255, 164, 128, or 64

OpenStudy (dtan5457):

It's 64. Even by logic from those choices..lol.

OpenStudy (dtan5457):

My equation has that at 64, so corret

OpenStudy (dtan5457):

*correct

OpenStudy (triciaal):

96-2(n-1) n=17 96 - 32=64